find equation of line passes through (5,-1)and is parallel to the line 2x+5y=15
Answer by Alan3354(69443)

find equation of line passes through (5,-1)and is parallel to the line 2x+5y=15
---------
Step 1, find the slope, m, of the line.
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is either point.